OpenStudy (anonymous):

The sum of two consecutive odd integers is 56. What is the smaller number? need answer /help

OpenStudy (anonymous):

Let's say you had a number - a random number, call it X.

OpenStudy (anonymous):

Do you know what a consecutive number is ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, what would a consecutive number to X look like ?

OpenStudy (radar):

Let x = a random integer, then we can say: 2x would be an even integer.......we could then say: 2x + 1 would be an odd integer. Now the next odd integer could be expressed as: 2x + 1 + 2 or 2x + 3 Now back to the problem:"The sum of two consecutive odd integers is 56. What is the smaller number? 2x + 1 first odd integer (and smallest) 2x + 3 2nd odd and consecutive integer. Now add ---------- 4x + 4 = 56 4x = 56 - 4 = 52 x = 13 2x + 1 = 27 First odd and smallest integer 2x + 3 = 29 Second and consecutive odd integer. 27 + 29 = 56 Verification.